# Finite Measures of Maximal Entropy for an Open Set of Partially Hyperbolic Diffeomorphisms

### v1

@@ 5. Proof of Theorems A and B
Fix $e^{h^{s}(f)-a}<\rho<1$ and consider $\Lambda_{\rho}\subset M$, and $\delta_{l}>0$ as in 2.10 and Remark 4.1 respectively.
@@
There exist $K_{a}\in\mathbb{N}$ and many finite $\mu_{1},\cdots\mu_{K_{a}}$ ergodic measures with $h_{\mu_{i}}(f)\geq a$ such that if $\mu\in{\mathcal{M}}_{f}^{e}$ with $h_{\mu}(f)\geq a$, then $\mu\in HC(\mu_{i})$ for some $i\leq K_{a}$.
@@
Now, let $\{\mu_{\gamma}\}$ be a sequence of different ergodic measures of $f$ such that $h_{\mu_{\gamma}}(f)\geq a$ and choose $x_{\mu_{\gamma}}$ holds 5.1. Due to the compactness of $M$, there are two measures $\mu_{\gamma_{1}}$ and $\mu_{\gamma_{2}}$ such that:
$d(x_{\mu_{\gamma_{1}}},x_{\mu_{\gamma_{2}}})<\frac{\delta_{l}}{3}.$
By Remark 4.1, for $\mu_{\gamma_{1}}$-almost every $x_{1}\in B_{\frac{\delta_{l}}{3}}(x_{\mu_{\gamma_{1}}})\cap\Lambda_{\rho}$ and $\mu_{\gamma_{2}}$-almost every $x_{2}\in B_{\frac{\delta_{l}}{3}}(x_{\mu_{\gamma_{2}}})\cap\Lambda_{\rho}$, we obtain that $W^{u/s}(x_{1})\pitchfork W^{s/u}(x_{2})$ are both not empty, and therefore, $\mu_{\gamma_{1}}\sim\mu_{\gamma_{2}}$.
Therefore, there is only a finite number of homoclinic classes in the set of hyperbolic measures containing measures with metric entropy greater than or equal to $a$. Finally, since every ergodic measure with metric entropy respect $f$ greater or equal than $a$ is a hyperbolic measure, we can choose $\mu_{1},\ldots,\mu_{K_{a}}$ as representatives of each distinct homoclinic class, each having entropy greater than or equal to $a$. Thus, we conclude the proof of the claim. ∎
To finish the proof, we simply choose $p_{1},\ldots,p_{K_{a}}$ with orbits ${{\mathcal{O}}}_{1},\ldots,{{\mathcal{O}}}_{K_{a}}$ such that ${{\mathcal{O}}}_{i}\sim\mu_{i}$ for every $1\leq i\leq K_{a}$ and $HC({\mathcal{O}}_{i})=HC_{T}(\mu_{i})$. This is possible by Proposition 2.7. ∎
@@
If $h^{u}(f)>h^{s}(f)$ then $f$ has a finite number of measures of maximal entropy.
@@
Finally, by Theorem A, the number of homoclinic classes in ${\mathcal{M}}_{f}^{h}$ containing a measure of maximal entropy is finite. Therefore $f$ has finite ergodic measures of maximal entropy.
@@
If $h^{u}(f)>h^{s}(f)$, then there exists a $C^{1}$ neighborhood ${\mathcal{U}}$ of $f$ such that every $C^{1+}$ diffeomorphism $g\in{\mathcal{U}}$ has a finite number of measures of maximum entropy.
@@
By Lemma 3.4 there exists a neighborhood ${\mathcal{U}}$ of $f$ such that if $g\in{\mathcal{U}}\cap\operatorname{Diff}^{1+}(M)$ so $h^{u}(g)>\frac{h^{u}(f)+h^{s}(f)}{2}>h^{s}(g)$. Finally, by Corollary 5.1 every $g\in{\mathcal{U}}\cap{{\rm Diff}}^{1+}(M)$ has many finite ergodic measures of maximal entropy.
@@
[…]
@@
If $\mu_{n}\in{\mathcal{M}}_{g_{n}}^{e}$ is a m.m.e entropy for $g_{n}$, then there exists an index $j$ and a subsequence $(\mu_{n_{l}})_{l\geq 0}$ such that $\mu_{n_{l}}\sim p_{j}(g_{n})$ for every $l\geq 0$.
@@
Let $\rho$ be a real number such that $h^{s}(g_{n})-h^{u}(g_{n})<\log(\rho)<0$ for every $n$ and let $\Lambda_{\rho}(g_{n})$ be as defined in 4.1. By Corollary 4.5, there exists $\beta>0$ such that $\Lambda_{\rho}(g_{n})\geq 2\beta$ and therefore we can choose a compact subset $\tilde{\Lambda}_{\rho}(g_{n})\subset\Lambda_{\rho}(g_{n})$ of regular points with $\mu(\tilde{\Lambda}_{\rho}(g_{n}))\geq\beta$. Let $n_{l}$ be an increased sequence of natural numbers such that $\mu_{n_{l}}$ converges to $\nu$ in the weak * topology and $\Lambda_{\rho}(g_{n_{l}})$ converges to a compact set $K$ in the Hausdorff metric. Observe that $K\subset\Lambda_{\rho}(f)$; so, by Remark 2.16, there is a compact subset $\tilde{\Lambda}_{\rho}(f)\subset K$ of regular points for $f$ with $\nu(\tilde{\Lambda}_{\rho}(f))\geq\frac{\beta}{2}$.
Observe that $\nu$ is not necessarily an ergodic measure however it is a maximal entropy measure of $f$. In effect, since $\mu_{n}\in{\mathcal{M}}_{g_{n}}^{e}$ and $\mu_{n}$ is m.m.e. for $g_{n}$, by Lemma 3.1, we have that (3.2) holds, that is
$h^{u}_{\mu_{n}}(g_{n})=h_{\mu_{n}}(g_{n})$
Furthermore, as shown in [33, Theorem A], we have:
@@
By Proposition 2.1 we can choose $x\in W^{s}(p_{j}(f))\cap W^{u}(p_{j}(f))$ such that
@@
$W^{s/u}_{\delta_{1}}(z)\pitchfork W^{s,u}_{\delta_{1}}(x)\neq\emptyset$
Since the invariant manifolds $W^{s/u}(p_{j}(g_{n_{l}}))$ converge to $W^{s/u}(p_{j}(f))$ in the $C^{1}$ topology, there exists a sequence $x_{n_{l}}\in W^{s}(p_{j}(g_{n_{l}}))\cap W^{u}(p_{j}(g_{n}))$ such that $(x_{n_{l}})_{{\mathbb{N}}}$ converges to $x$ and $W^{s/u}_{\delta_{1}}(x_{n_{l}})$ converges to $W^{s/u}_{\delta_{1}}(x)$ in the $C^{1}$ topology. Therefore, for $n$ large enough,
@@
Now, observe that $g_{n}$ cannot have two ergodic measures of maximal entropy that are homonically related to the same $p_{i}(g_{n})$ because they are not related to each other. Therefore, for sufficiently large $n$, every $g_{n}$ has at most $k$ ergodic measures of maximal entropy, otherwise, according to Lemma 6.3, there exists a subsequence $g_{n_{l}}$ such that every $g_{n_{l}}$ has at most two different ergodic measures of maximal entropy that are homoclinically related to the same $p_{i}(g_{n})$. Thus, the proof is concluded. ∎
6. Proof of Theorems C and D
In this section, we prove theorems C and D. As before, $f:M\to M$ denotes a $C^{1+}$ partially hyperbolic diffeomorphism of a compact manifold $M$ with $TM=E^{u}\oplus E^{c}\oplus E^{s}$ and ${\operatorname{dim}}(E^{c})=1$ and $\phi:M\to\mathbb{R}$ denote a potential Hölder continuous.
@@ 6. Proof of Theorems C and D
Finally, we choose the $K_{b}$ hyperbolic periodic orbits between $\mathcal{O}_{1},\cdots,\mathcal{O}_{K_{b}}$} which are homoclinically related to an ergodic measure that has metric pressure associated greater or equal to $b$. This concludes the proof. ∎
@@
If $h^{u}(f)>\sup\phi>\inf\phi>h^{s}(f)$ then $f$ has a finite number of measures of maximal entropy.
@@
As proven in Corollary 5.1. Subsection 2.3.1, Theorem 2.8, Proposition 2.7 and Lemma 3.3 imply that equilibrium states exist for $(f,\phi)$, each ergodic equilibrium state $\mu$ is hyperbolic and there exists an orbit ${{\mathcal{O}}}\in{{\rm Per}}_{h}(f)$ such that $\mu$ is the unique equilibrium state homoclinically related to it. Consequently, every homoclinic class in ${\mathcal{M}}_{f}^{h}$ can have, at most, one measure of maximal entropy.
By Theorem C the number of homoclinic classes in ${\mathcal{M}}_{f}^{h}$ with equilibrium states is finite. Therefore $(f,\phi)$ has a finite number of equilibrium states.
@@
If $h^{u}(f)>\sup\phi>\inf\phi>h^{s}(f)$, then there exists a $C^{1}$ neighborhood ${\mathcal{U}}$ of $f$ such that every $C^{1+}$ diffeomorphism $g\in{\mathcal{U}}$, $(g,\phi)$ has a finite number of equilibrium states.
@@
Taking $a_{1}=\sup\phi$ and $a_{2}=\inf\phi$ in Lemma 3.4 there exists a $C^{1}$ neighborhood ${\mathcal{U}}$ of $f$ such that for every $g\in{{\rm Diff}}^{1+}(M)\cap{\mathcal{U}}$ holds $h^{u}(g)>\sup\phi>\inf\phi>h^{s}(g)$. So, applying Corollary 6.1 we get that each $g\in{{\rm Diff}}^{1+}(M)\cap{\mathcal{U}}$ is such that $(g,\phi)$ has a finite number of ergodic equilibrium states. ∎
Proof of Theorem D.
Suppose that $h^{u}(f)>\sup\phi>\inf\phi>h^{s}(f)$. Consider $a_{1}=\sup\phi$ and $a_{2}=\inf\phi$, by Lemma 3.4 and Corollary 6.2 there exists a $C^{1}$ neighborhood ${\mathcal{U}}$ of $f$ such that each $g\in{{\rm Diff}}^{1+}(M)\cap{\mathcal{U}}$ holds $h^{u}(f)>\sup\phi>\inf\phi>h^{s}(f)$ and $(g,\phi)$ has many finite ergodic equilibrium states.
By the Corollary 6.1 the number of equilibrium estates for $(f,\phi)$ is $k>0$. Let $\nu_{1},\cdots\nu_{k}$ be these ergodic equilibrium states and $p_{1}(f)\cdots p_{k}(f)$ the periodic hyperbolic points with orbits ${{\mathcal{O}}}_{1}\cdots{\O}_{k}$ respectively of $f$ such that ${{\mathcal{O}}}_{i}\sim\mu_{i}$ and
$P(f_{|HC(\mathcal{O}_{i})},\phi_{|HC(\mathcal{O}_{i})})=P(f,\phi)$
@@
If $\mu_{n}\in{\mathcal{M}}_{g_{n}}^{e}$ is an equilibrium state for $g_{n}$, then there exists an index $j$ and a subsequence $(\mu_{n_{l}})_{l\geq 0}$ such that $\mu_{n_{l}}\sim p_{j}(g_{n})$ for every $l\geq 0$.
@@
We will proceed as in Lemma 6.3. Let $\rho$ be a real number as in Corollary 4.7 and for every $n$ consider $\Lambda_{\rho}(g_{n})$ as in 4.1. By Corollary 4.7, there exists $\beta>0$ such that $\Lambda_{\rho}(g_{n})\geq 2\beta$ and therefore we can choose a compact subset $\tilde{\Lambda}_{\rho}(g_{n})\subset\Lambda_{\rho}(g_{n})$ of regular points with $\mu(\tilde{\Lambda}_{\rho}(g_{n}))\geq\beta$. Let $n_{l}$ be an increased sequence of natural numbers such that $\mu_{n_{l}}$ converges to $\nu$ in the weak * topology and $\Lambda_{\rho}(g_{n_{l}})$ converges to a compact set $K$ in the Hausdorff metric. Observe that $K\subset\Lambda_{\rho}(f)$; so, by Remark 2.16, there is a subset $\tilde{\Lambda}_{\rho}(f)\subset K$ as a compact set of regular points for $f$ with $\nu(\tilde{\Lambda}_{\rho}(f))\geq\frac{\beta}{2}$.
The measure $\nu$ is an equilibrium state of $(f,\phi)$. In effect, since $\mu_{n}\in{\mathcal{M}}_{g_{n}}^{e}$ and $\mu_{n}$ is an equilibrium state of $g_{n}$, by Lemma 3.3, we have that
$h^{u}_{\mu_{n}}(g_{n})=h_{\mu_{n}}(g_{n})$
Furthermore, as shown in [33, Theorem A], we have:
@@
Therefore, since each $\mu_{n}$ is an equilibrium state of $g_{n}$, the topological pressure is a continuous function on ${{\mathcal{U}}}\cap{{\rm Diff}}^{r+1}(M)$ and $h^{u}_{\mu}(f)\leq h_{\mu}(f)$, we conclude that $\nu$ is also an equilibrium state of $(f,\phi)$.
Thus, the measure $\nu$ is a convex combination of $\nu_{1},\cdots,\nu_{k}$, and there is $j$ such that $\nu_{j}(\tilde{\Lambda}_{\rho}(f))\geq\frac{\beta}{2}$. Consider $p_{j}(f)$ such that $p_{j}\sim\mu_{j}$ and repeating the argument in Lemma 6.3, by the continuity of the invariant manifolds, we conclude that for every $l$ large enough $p_{i}(g_{n_{l}})\sim\mu_{n_{l}}$. Thus, the proof is concluded. ∎
Finally, with the same argument of the proof of Theorem B, we conclude $g_{n}$ has at most $k$ ergodic equilibrium states when $n$ is big enough.
Acknowledgements. We are thankful to Jiagang Yang for helpful conversations on this work.

### v3

@@ 5. Proof of Theorems A and B
Fix $e^{h^{s}(f)-a}<\rho<1$ and consider $\Lambda_{\rho}\subset M$, and $\delta_{l}>0$ as in Definition 2.10 and Remark 4.1 respectively.
@@
There exist $K_{a}\in\mathbb{N}$ and finitely many ergodic measures $\mu_{1},\cdots\mu_{K_{a}}$ with $h_{\mu_{i}}(f)\geq a$ such that if $\mu\in{\mathcal{M}}_{f}^{e}$ with $h_{\mu}(f)\geq a$, then $\mu\in HC(\mu_{i})$ for some $i\leq K_{a}$.
@@
Now, let $(\mu_{j})_{j\in\mathbb{N}}$ be a sequence of distinct ergodic measures of $f$ such that $h_{\mu_{j}}(f)\geq a$ for every ${j\in\mathbb{N}}$ and choose $x_{\mu_{j}}$ such that (5.1) holds. Due to the compactness of $M$, there are $n,\,m\,\in\mathbb{N},n\neq m,$ such that:
$d(x_{\mu_{n}},x_{\mu_{m}})<\frac{\delta_{l}}{3}.$
By Remark 4.1, for $\mu_{n}$-almost every $x_{1}\in B_{\frac{\delta_{l}}{3}}(x_{\mu_{n}})\cap\Lambda_{\rho}$ and $\mu_{m}$-almost every $x_{2}\in B_{\frac{\delta_{l}}{3}}(x_{\mu_{m}})\cap\Lambda_{\rho}$, we obtain that $W^{u/s}(x_{1})\pitchfork W^{s/u}(x_{2})$ are both not empty, and therefore, $\mu_{n}\sim\mu_{m}$.
Therefore, there is only a finite number of homoclinic classes in the set of hyperbolic measures containing measures with metric entropy greater than or equal to $a$. Finally, since every ergodic measure with metric entropy greater or equal than $a$ respect $f$ is a hyperbolic measure, we can choose $\mu_{1},\ldots,\mu_{K_{a}}$ as representatives of each distinct homoclinic class, each having entropy greater than or equal to $a$. Thus, we conclude the proof of the claim. ∎
To finish the proof, we simply choose $p_{1},\ldots,p_{K_{a}}\in M$ with hyperbolic periodic orbits ${{\mathcal{O}}}_{1},\ldots,{{\mathcal{O}}}_{K_{a}}$ such that ${{\mathcal{O}}}_{i}\sim\mu_{i}$ for every $1\leq i\leq K_{a}$ and $HC({\mathcal{O}}_{i})=HC_{T}(\mu_{i})$. This is possible by Proposition 2.7. ∎
@@
If $h^{u}(f)>h^{s}(f)$ then $f$ has only a finite number of measures of maximal entropy.
@@
Finally, by Theorem A, the number of homoclinic classes in ${\mathcal{M}}_{f}^{h}$ containing a measure of maximal entropy is finite. Therefore $f$ has finitely many ergodic measures of maximal entropy.
@@
If $h^{u}(f)>h^{s}(f)$, then there exists a $C^{1}$ neighborhood ${\mathcal{U}}$ of $f$ such that every $C^{1+}$ diffeomorphism $g\in{\mathcal{U}}$ has only a finite number of measures of maximum entropy.
@@
By Lemma 3.4 there exists a neighborhood ${\mathcal{U}}$ of $f$ such that if $g\in{\mathcal{U}}\cap\operatorname{Diff}^{1+}(M)$ so $h^{u}(g)>\frac{h^{u}(f)+h^{s}(f)}{2}>h^{s}(g)$. Finally, by Corollary 5.1 every $g\in{\mathcal{U}}\cap{{\rm Diff}}^{1+}(M)$ has finitely many ergodic measures of maximal entropy.
@@
Let us consider a sequence of $C^{1+}$ diffeomorphisms, denoted as $g_{n}$, where each $g_{n}$ belongs to ${\mathcal{U}}$ and the sequence $(g_{n})_{n}$ converges to $f$ in the $C^{1}$ topology. Let $p_{i}(g_{n})$ be the unique hyperbolic periodic point of $g_{n}$ derived from the structural stability of $p_{i}(f)$.
@@
If $\mu_{n}\in{\mathcal{M}}_{g_{n}}^{e}$ is a m.m.e. for $g_{n}$, then there exists an index $j$ and a subsequence $(\mu_{n_{l}})_{l\geq 0}$ such that $\mu_{n_{l}}\sim p_{j}(g_{n_{l}})$ for every $l\geq 0$.
@@
Let $\rho$ be a real number such that $h^{s}(g_{n})-h^{u}(g_{n})<\log(\rho)<0$ for every $n$ and let $\Lambda_{\rho}(g_{n})$ be as defined in (4.1). By Corollary 4.5, there exists $\beta>0$ such that $\mu_{n}(\Lambda_{\rho}(g_{n}))\geq 2\beta$ and therefore we can choose a compact subset $\tilde{\Lambda}_{\rho}(g_{n})\subset\Lambda_{\rho}(g_{n})$ of regular points of $\mu_{n}$ with $\mu_{n}(\tilde{\Lambda}_{\rho}(g_{n}))\geq\beta$. Let $n_{l}$ be an increased sequence of natural numbers such that $\mu_{n_{l}}$ converges to $\nu$ in the weak * topology and $\Lambda_{\rho}(g_{n_{l}})$ converges to a compact set $K$ in the Hausdorff metric. Observe that $K\subset\Lambda_{\rho}(f)$; so, by Remark 2.16, there is a compact subset $\tilde{\Lambda}_{\rho}(f)\subset K$ of regular points for $f$ with $\nu(\tilde{\Lambda}_{\rho}(f))\geq\frac{\beta}{2}$.
Observe that $\nu$ is not necessarily an ergodic measure however it is a maximal entropy measure of $f$. In fact, since $\mu_{n}\in{\mathcal{M}}_{g_{n}}^{e}$ and $\mu_{n}$ is m.m.e. for $g_{n}$, by Lemma 3.1, we have that (3.2) holds, that is,
$h^{u}_{\mu_{n}}(g_{n})=h_{\mu_{n}}(g_{n}).$
Furthermore, as shown in [53, Theorem A], we have:
@@
By Definition 2.1 we can choose $x\in W^{s}(p_{j}(f))\cap W^{u}(p_{j}(f))$ such that
@@
$W^{s/u}_{\delta_{1}}(z)\pitchfork W^{s,u}_{\delta_{1}}(x)\neq\emptyset.$
Since the invariant manifolds vary continuously with $g_{n}$, the invariant manifolds $W^{s/u}(p_{j}(g_{n_{l}}))$ converge to $W^{s/u}(p_{j}(f))$ in the $C^{1}$ topology, and so there exists a sequence $x_{n_{l}}\in W^{s}(p_{j}(g_{n_{l}}))\cap W^{u}(p_{j}(g_{n}))$ such that $(x_{n_{l}})_{l\in{\mathbb{N}}}$ converges to $x$ and $W^{s/u}_{\delta_{1}}(x_{n_{l}})$ converges to $W^{s/u}_{\delta_{1}}(x)$ in the $C^{1}$ topology. Therefore, for $n$ large enough,
@@
Now, observe that $g_{n}$ cannot have two ergodic measures of maximal entropy that are homonically related to the same $p_{i}(g_{n})$. Thus, the proof is concluded. ∎
Let $\mu_{1},\cdots,\mu_{k}$ be the ergodic measures of maximal entropy of $f$. Also, let $p_{i}(f)$ denote the hyperbolic point associated to $\mu_{i}$ for $i=1,\cdots k$. For a $C^{1}$ diffeomorphism $g$ sufficiently close to $f$, consider the hyperbolic periodic points $p_{i}(g)$ such that each $p_{i}(g)$ is near $p_{i}(f),\,i=1,\cdots k$. We obtain the following corollary:
Corollary 5.4.
Let $f:M\to M$ be a $C^{1+}$ partially hyperbolic diffeomorphism of a closed manifold $M$ with $TM=E^{u}\oplus E^{c}\oplus E^{s}$ and $\operatorname{dim}(E^{c})=1$. If $h^{u}(f)>h^{s}(f)$, then there exists a $C^{1}$ neighborhood $\mathcal{U}$ of $f$ such that for every $C^{1+}$ diffeomorphism $g\in\mathcal{U}$, if $\mu$ is an ergodic measure of maximal entropy for $g$ then $\mu$ is homoclinically related to the orbit of $p_{j}(g)$ for some $j\in{1,\cdots,k}$.
Proof.
Assume, by contradiction, the existence of a sequence of $C^{1+}$ diffeomorphisms $(g_{n})_{n\in\mathbb{N}}$ converging to $f$, and let $\mu_{n}$ be an ergodic measure of maximal entropy of $g_{n}$ such that $\mu_{n}$ is not homoclinically related to any $p_{i}(g_{n})$ for $i=1,\cdots,k$. This sequence contradicts Lemma 5.3. Thus, we conclude the proof. ∎
6. Proof of theorem C
In this section, we prove Theorem C. Recall that $f:M\to M$ denotes a $C^{1+}$ partially hyperbolic diffeomorphism of a compact manifold $M$ with $TM=E^{u}\oplus E^{c}\oplus E^{s}$ and ${\rm dim}(E^{c})=1$. Moreover, the sets ${\mathcal{M}}^{+}_{f}$ and ${\mathcal{M}}^{-}_{f}$ are defined as follows:
${\mathcal{M}}^{+}_{f}=\{\mu\in{\mathcal{M}}_{f}^{e}:\lambda_{c}(f,\mu)\geq 0\}\mbox{ and }{\mathcal{M}}^{-}_{f}=\{\mu\in{\mathcal{M}}_{f}^{e}:\lambda_{c}(f,\mu)\leq 0\}$
and the values $h^{-}(f)$ and $h^{-}(f)$ are given by
$h^{-}(f)=\sup_{\mu\in{\mathcal{M}}_{f}^{-}}\{h_{\mu}(f)\}\mbox{ and }h^{+}(f)=\sup_{\mu\in{\mathcal{M}}_{f}^{+}}\{h_{\mu}(f)\}.$
Note that if $h^{+}(f)\neq h^{-}(f)$, then either $h^{+}(f)>h^{-}(f)$ or $h^{+}(f)<h^{-}(f)$. We will consider the case where $h^{-}(f)>h^{+}(f)$; the proof for $h^{+}(f)>h^{-}(f)$ is analogous. We prove the following proposition.
Proposition 6.1.
Let $f:M\to M$ be a $C^{1+}$ partially hyperbolic diffeomorphism of a compact manifold $M$ with $TM=E^{u}\oplus E^{c}\oplus E^{s}$ and $\operatorname{dim}(E^{c})=1$. Then, $h^{-}(f)>h^{+}(f)$ if and only if there exist $\lambda<0$ and $h(f)\geq\tilde{h}>0$ such that if $\mu$ is an ergodic measure of $f$ with $h_{\mu}(f)>\tilde{h}$, then $\lambda_{c}(f,\mu)<-\lambda$.
Proof.
Sufficiency has been proven in [12, Equation (9.2)], and necessity follows directly from the definition of $h^{-}$ and $h^{+}$. ∎
Lemma 6.2.
Suppose that $h^{-}(f)>h^{+}(f)$. If $\mu\in{\mathcal{M}}_{f}^{e}$ with $h_{\mu}(f)>\tilde{h}$ then for every $e^{-\lambda}<\rho<1$ we have that $\mu(\Lambda_{\rho})\geq\frac{\log(\rho)-\lambda_{c}(f,\mu)}{\log(\rho)-m(f)}.$
The proof of this lemma is analogous to the proof of Lemma 4.3 and so we omit it.
@@ 6. Proof of theorem C
Note that for $a>\tilde{h}$ by Lemma 6.2 we have the same conclusion than Theorem A. So, we can conclude that $f$ has finitely many ergodic measure of maximal entropy.
7. Proof of Theorems D and E
In this section, we prove Theorems D and E. As before, $f:M\to M$ denotes a $C^{1+}$ partially hyperbolic diffeomorphism of a compact manifold $M$ with $TM=E^{u}\oplus E^{c}\oplus E^{s}$ and ${\operatorname{dim}}(E^{c})=1$ and $\phi:M\to\mathbb{R}$ denotes a Hölder continuous potential.
Proof of Theorem D.
@@ 7. Proof of Theorems D and E
Finally, we choose the $K_{b}$ hyperbolic periodic orbits $\mathcal{O}_{1},\cdots,\mathcal{O}_{K_{b}}$} which are homoclinically related to an ergodic measure that has metric pressure associated greater or equal to $b$. This concludes the proof. ∎
@@
If $h^{u}(f)>\sup\phi>\inf\phi>h^{s}(f)$ then $f$ has only a finite number of ergodic equilibirum states.
@@
As proven in Corollary 5.1. Subsection 2.3.1, Theorem 2.8, Proposition 2.7 and Lemma 3.3 imply that equilibrium states exist for $(f,\phi)$, each ergodic equilibrium state $\mu$ is hyperbolic and there exists an orbit ${{\mathcal{O}}}\in{{\rm Per}}_{h}(f)$ such that $\mu$ is the unique equilibrium state homoclinically related to it. Consequently, every homoclinic class in ${\mathcal{M}}_{f}^{h}$ can have, at most, one ergodic equilibirum states.
By Theorem D the number of homoclinic classes in ${\mathcal{M}}_{f}^{h}$ with equilibrium states is finite. Therefore $(f,\phi)$ has only a finite number of ergodic equilibrium states.
@@
If $h^{u}(f)>\sup\phi>\inf\phi>h^{s}(f)$, then there exists a $C^{1}$ neighborhood ${\mathcal{U}}$ of $f$ such that for every $C^{1+}$ diffeomorphism $g\in{\mathcal{U}}$, $(g,\phi)$ has only a finite number of equilibrium states.
@@
Taking $a_{1}=\sup\phi$ and $a_{2}=\inf\phi$ in Lemma 3.4 there exists a $C^{1}$ neighborhood ${\mathcal{U}}$ of $f$ such that for every $g\in{{\rm Diff}}^{1+}(M)\cap{\mathcal{U}}$ it holds $h^{u}(g)>\sup\phi>\inf\phi>h^{s}(g)$. So, applying Corollary 7.1 we get that each $g\in{{\rm Diff}}^{1+}(M)\cap{\mathcal{U}}$ is such that $(g,\phi)$ has only a finite number of ergodic equilibrium states. ∎
Proof of Theorem E Suppose that $h^{u}(f)>\sup\phi>\inf\phi>h^{s}(f)$. Consider $a_{1}=\sup\phi$ and $a_{2}=\inf\phi$, by Lemma 3.4 and Corollary 7.2 there exists a $C^{1}$ neighborhood ${\mathcal{U}}$ of $f$ such that each $g\in{{\rm Diff}}^{1+}(M)\cap{\mathcal{U}}$ holds $h^{u}(f)>\sup\phi>\inf\phi>h^{s}(f)$ and $(g,\phi)$ has finitely many ergodic equilibrium states.
By Corollary 7.1 the number of equilibrium estates for $(f,\phi)$ is $k>0$. Let $\nu_{1},\cdots\nu_{k}$ be these ergodic equilibrium states and $p_{1}(f)\cdots p_{k}(f)$ the periodic hyperbolic points with orbits ${\mathcal{O}}_{1}\cdots{{\mathcal{O}}}_{k}$ respectively of $f$ such that ${{\mathcal{O}}}_{i}\sim\mu_{i}$ and
$P(f_{|HC(\mathcal{O}_{i})},\phi_{|HC(\mathcal{O}_{i})})=P(f,\phi).$
@@
If $\mu_{n}\in{\mathcal{M}}_{g_{n}}^{e}$ is an equilibrium state for $g_{n}$, then there exists an index $j$ and a subsequence $(\mu_{n_{l}})_{l\geq 0}$ such that $\mu_{n_{l}}\sim p_{j}(g_{n_{l}})$ for every $l\geq 0$.
@@
We will proceed as in Lemma 5.3. Let $\rho$ be a real number as in Corollary 4.6 and for every $n$ consider $\Lambda_{\rho}(g_{n})$ as in (4.1). By Corollary 4.6, there exists $\beta>0$ such that $\Lambda_{\rho}(g_{n})\geq 2\beta$ and therefore we can choose a compact subset $\tilde{\Lambda}_{\rho}(g_{n})\subset\Lambda_{\rho}(g_{n})$ of regular points with $\mu(\tilde{\Lambda}_{\rho}(g_{n}))\geq\beta$. Let $n_{l}$ be an increasing sequence of natural numbers such that $\mu_{n_{l}}$ converges to $\nu$ in the weak * topology and $\Lambda_{\rho}(g_{n_{l}})$ converges to a compact set $K$ in the Hausdorff metric. Observe that $K\subset\Lambda_{\rho}(f)$; so, by Remark 2.16, there is a subset $\tilde{\Lambda}_{\rho}(f)\subset K$ as a compact set of regular points for $f$ with $\nu(\tilde{\Lambda}_{\rho}(f))\geq\frac{\beta}{2}$.
The measure $\nu$ is an equilibrium state of $(f,\phi)$. In effect, since $\mu_{n}\in{\mathcal{M}}_{g_{n}}^{e}$ is an equilibrium state of $(g_{n},\phi)$, by Corollary 3.3 and Equation (3.1), we have that
$h^{u}_{\mu_{n}}(g_{n})=h_{\mu_{n}}(g_{n}).$
Furthermore, as shown in [53, Theorem A], we have:
@@
Therefore, since each $\mu_{n}$ is an equilibrium state for $(g_{n},\phi)$, the topological pressure is a continuous function on ${{\mathcal{U}}}\cap{{\rm Diff}}^{r+1}(M)$ , and $h^{u}_{\mu}(f)\leq h_{\mu}(f)$, we conclude that $\nu$ is also an equilibrium state of $(f,\phi)$.
Thus, the measure $\nu$ is a convex combination of $\nu_{1},\cdots,\nu_{k}$, and there is $j$ such that $\nu_{j}(\tilde{\Lambda}_{\rho}(f))\geq\frac{\beta}{2}$. Consider $p_{j}(f)$ such that $p_{j}\sim\mu_{j}$ and repeating the argument in 7.3 Theorem B , by the continuity of the invariant manifolds with $g_{n}$, we conclude that for every $l$ large enough $p_{i}(g_{n_{l}})\sim\mu_{n_{l}}$. Thus, the proof is concluded.
Finally, with the same argument of the proof of Theorem B, we conclude that $g_{n}$ has at most $k$ ergodic equilibrium states when $n$ is big enough.
Acknowledgements. We are thankful to Jiagang Yang and Fan Yang for helpful conversations on this work.